# Khatri-Rao Clustering for Data Summarization

### v1

@@ 4. Problem Formulation
Although Khatri-Rao clustering defines a general paradigm for centroid-based clustering, we focus on $k$-Means and two deep clustering algorithms. Khatri-Rao extensions of other centroid-based clustering algorithms, e.g., based on matrix decomposition (Li and Ding, 2013) or gradient descent (Sculley, 2010), are possible but require method-specific adjustments. Investigating whether our ideas can also be applied to clustering algorithms that do not rely on centroids, such as hierarchical clustering (Ghosal et al., 2020), is left to future research.
@@ 4.1. Khatri-Rao $k$-Means Clustering
The classic $k$-Means problem is known to be $\mathbf{NP}$-Hard (Dasgupta, 2008). Any instance of $k$-Means can be mapped to an instance of Khatri-Rao $k$-Means simply by setting $p=1$ and $h_{1}=k$. Thus, Problem 1 is also $\mathbf{NP}$-hard.
@@ 4.2. Khatri-Rao Deep Clustering
In the autoencoder-based deep clustering setting, autoencoders consist of an encoder $f^{e}_{\Theta_{\boldsymbol{\alpha}}}:\mathcal{D}\to Z$ mapping data to latent representations $Z=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, and a decoder $f^{d}_{\Theta_{\boldsymbol{\alpha}}}:Z\to\hat{\mathcal{D}}$, reconstructing data from their latent representations. We consider an autoencoder with $n_{l}$ layers. The $l$-th layer is parameterized by a matrix $\mathbf{W}_{l}\in\mathbb{R}^{d_{l}\times m_{l}}$, which could be relatively large. To obtain a succinct representation of the output of deep clustering algorithms, it is thus not sufficient to compress the centroid parameters, but it is also necessary to compress the autoencoder parameters $\Theta_{\boldsymbol{\alpha}}=\{\mathbf{W}_{l}\}_{l=1}^{n_{l}}$. To this aim, we draw on the connection between Khatri-Rao clustering and Hadamard decomposition (Ciaperoni et al., 2024) discussed in Section 2, and we reparametrize matrices $\mathbf{W}_{l}$ as Hadamard products of $q$ factors, i.e.:
@@
where $\mathbf{A}^{l}_{i}\in\mathbb{R}^{d_{l}\times r_{i}}$ and $\mathbf{B}^{l}_{i}\in\mathbb{R}^{r_{i}\times m_{l}}$. The key intuition behind such reparameterization mirrors that of Khatri-Rao clustering. The Hadamard product of $q$ matrices of ranks $r_{1},r_{2},\dots$ and $r_{q}$ can represent matrices of rank up to $\prod_{l=1}^{q}r_{l}$ while using only $2\sum_{l=1}^{q}r_{l}$ vectors. In practice, this provides an effective compression mechanism, as also demonstrated in the context of computer vision in federated learning environments (Hyeon-Woo et al., 2021), where a similar reparameterization is considered.
@@ 8. Design Choices in Khatri-Rao Clustering
Choosing the number of centroids. If the budget $b$ of vectors is not fixed and one wants to identify the number of clusters that are appropriate for the data, Khatri-Rao clustering can be combined with established techniques such as $X$-Means (Pelleg and Moore, 2000) or $G$-Means (Hamerly and Elkan, 2003). Here, the number of centroids is successively increased and the current parameterization is evaluated, e.g., by using the Bayesian Information Criterion (Schwarz, 1978) or by testing if certain distributional conditions are fulfilled. In Khatri-Rao clustering, increasing the number of clusters is equivalent to either increasing the cardinality of one set of protocentroids or the number of sets of protocentroids. Moreover, for deep clustering, a Khatri-Rao variant of such algorithms as DipDECK (Leiber et al., 2021) or DeepDPM (Ronen et al., 2022), which also optimize the number of underlying clusters, can be designed.
@@ 9.1. Settings
Datasets. We consider two synthetic dataset generators (Blobs and Classification) from the scikit-learn Python library (Pedregosa et al., 2011), which allow exploring the behavior of different algorithms as a function of the number of data points, features and underlying clusters. The preliminary experiments and the scalability analysis focus on those datasets.
In addition, we consider seven real-world benchmark datasets, available through the ClustPy (Leiber et al., 2023) Python library and two synthetic datasets available through the clustbench (Gagolewski, 2022) Python library. Finally, we consider the Double MNIST dataset, which is created by concatenating pairs of digits from MNIST (LeCun et al., 1998), yielding $100$ clusters representing the numbers 0 to $99$.
@@
In all deep clustering experiments, the parameter $w_{rec}$, which balances the clustering and reconstruction losses as per Equation (2), is set to $1$, the batch size for batch-wise gradient-based optimization is set to $512$, and we use the ADAM optimizer (Kingma and Ba, 2015) using a learning rate of $10^{-3}$ for pre-training the autoencoder and $10^{-4}$ for clustering.
@@
Metrics. In our experiments, the accuracy of a data summary is measured by the quality of clustering results. We consider the following standard metrics to evaluate clustering results that take advantage of ground truth labels: adjusted Rand index (ARI) (Hubert and Arabie, 1985), unsupervised clustering accuracy (ACC) (Yang et al., 2010) and normalized mutual information (NMI) (Kvalseth, 1987). These metrics have a maximum value of $1$, which reveals a perfect match between predicted and ground-truth labels. In the case of $k$-Means clustering, we also monitor inertia, which is the objective function optimized for by standard $k$-Means and Khatri-Rao-$k$-Means. In addition, in the preliminary experiments, we monitor clustering purity, which measures the proportion of correctly assigned data points after assigning each data point to the majority ground-truth label of its cluster (Manning et al., 2008).
@@
Clustering algorithms. In the case of $k$-Means clustering, we compare Khatri-Rao-$k$-Means against the standard $k$-Means algorithm (implementation from scikit-learn (Pedregosa et al., 2011)). Additionally, we consider the two-phase naïve approach described in Section 5 with product aggregator. When not clear from context, we specify the aggregator function and the budget of vectors used to represent centroids (e.g., Khatri-Rao-$k$-Means-++ $(h_{1}+h_{2})$ denotes Khatri-Rao-$k$-Means with sum aggregator and $h_{1}+h_{2}$ protocentroids).
In the case of deep clustering, we consider DKM and IDEC (implementations from ClustPy (Leiber et al., 2023)) against their Khatri-Rao variants, Khatri-Rao DKM and Khatri-Rao IDEC. Appendix B provides implementation details for all algorithms.
@@ 9.4. Case Studies
Obtaining more succinct codebooks for color quantization. Color quantization in computer graphics enables efficient compression of certain types of images by reducing the number of colors. It allows displaying images with many colors on hardware-constrained devices that can only display a limited number of colors, usually due to memory limitations (Brun and Trémeau, 2017). A common approach casts color quantization as clustering in a three-dimensional RGB space, where pixels are points. The centroids obtained by $k$-Means form a codebook of representative colors, and each pixel is mapped to its closest centroid for color quantization.
@@
Reducing communication costs for clustering in federated learning environments. In recent years, federated learning, which allows for collaborative training of machine learning models across distributed entities, has emerged as a central area of research (Wen et al., 2023). The Khatri-Rao clustering paradigm can benefit federated learning by reducing communication costs between entities, an important goal in federated learning research (Paragliola and Coronato, 2022).
Recently, Garst and Reinders (2024) have introduced an effective implementation of $k$-Means in a federated setting (henceforth called FkkM). As FkkM requires communicating centroids between several clients and a server, Khatri-Rao-FkkM can potentially reduce the costs of communication by considering a lightweight set of protocentroids instead of regular centroids. To extend FkkM to Khatri-Rao clustering, it suffices to replace each invocation of $k$-Means in FkkM with Khatri-Rao-$k$-Means. The resulting algorithm is referred to as Khatri-Rao-FkkM. To demonstrate the advantages given by Khatri-Rao-FkkM with respect to communication costs, we simulate a federated learning environment with $10$ clients and one server. In this experiment, we consider the benchmark FEMNIST dataset (Caldas et al., 2018), and we measure the inertia achieved by FkkM and Khatri-Rao-FkkM as a function of the communication costs from the server to the clients. The results of this experiment are displayed in Figure 10 and indicate that Khatri-Rao-FkkM consistently reduces inertia relative to FkkM at parity communication cost. For the smallest communication cost, the inertia incurred by FkkM is about five times larger than that incurred by Khatri-Rao-FkkM. This suggests that Khatri-Rao-FkkM can communicate clustering results of a given quality much more efficiently than FkkM.
@@ Appendix A Dataset Details
  • The MNIST dataset (LeCun et al., 1998), available via PyTorch (Paszke et al., 2017) or ClustPy (Leiber et al., 2023), is a collection of $28\times 28$ (vectorized) images of handwritten grayscale digits. We draw a stratified subsample of $25000$ images, ensuring that the original class proportions across the ten digit clusters are preserved. We rescale all pixel values by dividing by the maximum.
@@
  • The HAR dataset, available via the UCI repository (Kelly et al., 2023) or ClustPy (Leiber et al., 2023), consists of sensor readings collected from smartphone accelerometers and gyroscopes during human activity monitoring. There are $10299$ data points. Each data point is represented as a multivariate feature vector of dimension $561$ derived from raw time-series measurements, and the dataset includes $6$ activity clusters such as walking, standing and sitting. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Olivetti Faces dataset, available via ClustPy (Leiber et al., 2023) or scikit-learn (Pedregosa et al., 2011) contains (vectorized) grayscale facial images of $40$ individuals, with $10$ images per subject captured under varying lighting conditions, facial expressions, and poses. Each image has resolution $64\times 64$ pixels. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The CMU Faces dataset, available via the UCI repository (Kelly et al., 2023) or ClustPy (Leiber et al., 2023), contains $624$ grayscale facial images of $20$ persons varying their pose (up, straight, left and right) and expression (neutral, happy, sad, angry), and shown with and without sunglasses. The original images with a resolution of $30\times 32$ pixels are vectorized and each feature is standardized by subtracting its mean and dividing by its standard deviation.
@@
  • The Symbols dataset, available via ClustPy [^6] The dataset is also available at: https://www.timeseriesclassification.com/index.php (Leiber et al., 2023), consists of vectorized handwritten symbols. Each sample corresponds to a time series obtained from the drawing trajectory of a symbol. The symbols are drawn by $13$ different individuals. The dataset contains a total of $1020$ data points, each with $398$ measurements. The data points are organized into $6$ natural clusters. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The stickfigures dataset (Günnemann et al., 2014), available via ClustPy (Leiber et al., 2023), consists of $900$ synthetic (vectorized) silhouette images of human stick figures generated under varying poses. Each image has a resolution of $20\times 20$ pixels. Examples of such images are given in Figure 1. Each image is provided at a fixed resolution and captures a simplified body configuration defined by joint positions and limb orientations, resulting in $9$ distinct pose-based clusters. We rescale all pixel values by dividing by the maximum.
@@
  • The optdigits dataset, available via the UCI repository (Kelly et al., 2023) or ClustPy (Leiber et al., 2023), contains $5620$ (vectorized) handwritten digit images represented as $8\times 8$ grayscale pixel grids. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Classification dataset, sourced by scikit-learn (Pedregosa et al., 2011), is a synthetic dataset. While it was originally designed for benchmarking classifiers, it is also useful for evaluating clustering algorithms which simply discard class labels until the evaluation of the clustering results. Each class for classification corresponds to a cluster. Except in the experiments where we vary the number of data points, features, or clusters, the dataset consists of $5000$ samples organized into $100$ clusters. The data are generated with $10$ informative features, with no redundant or repeated features, ensuring that all dimensions contribute meaningfully to class separability. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Chameleon dataset, available via the clustbench Benchmark Suite (Gagolewski, 2022), consists of $10000$ two-dimensional point clouds exhibiting complex, nonconvex cluster shapes with varying densities (Karypis et al., 1999). The dataset corresponds to the configuration shown at the bottom of Figure 4, augmented with a substantial proportion of uniformly distributed data points that do not belong to any natural cluster.
@@
  • The Soybean Large dataset, available via the UCI repository (Kelly et al., 2023) or ClustPy (Leiber et al., 2023), is a dataset of categorical features. It consists of $562$ plant samples belonging to one of $15$ classes. For each plant, there are $35$ observed categorical attributes describing the plant. After encoding categorical attributes numerically, we standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Blobs dataset, sourced by scikit-learn (Pedregosa et al., 2011), is a synthetic dataset specifically designed for controlled evaluation of clustering algorithms that consists of data points grouped around isotropic Gaussian clusters. Each cluster has standard deviation $1$. The dataset consists of $5000$ 22-dimensional data points, arranged in $100$ clusters, except in the experiments where we vary the number of data points, features and clusters. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The R15 dataset, available via the clustbench benchmark suite (Gagolewski, 2022), consists of $600$ two-dimensional data points forming $15$ Gaussian clusters. The cluster centroids in R15 are not arranged on a regular grid but exhibit non-uniform inter-cluster distances. We standardize each feature by subtracting its mean and dividing by its standard deviation.

### v2

@@ 4. Problem Formulation
Although Khatri-Rao clustering defines a general paradigm for centroid-based clustering, we focus on $k$-Means and two deep clustering algorithms. Khatri-Rao extensions of other centroid-based clustering algorithms, e.g., based on matrix decomposition (li2013nonnegative) or gradient descent (DBLP:conf/www/Sculley10), are possible but require method-specific adjustments. Investigating whether our ideas can also be applied to clustering algorithms that do not rely on centroids, such as hierarchical clustering (ghosal2020short), is left to future research.
@@ 4.1. Khatri-Rao $k$-Means Clustering
The classic $k$-Means problem is known to be $\mathbf{NP}$-Hard (dasgupta2008hardness). Any instance of $k$-Means can be mapped to an instance of Khatri-Rao $k$-Means simply by setting $p=1$ and $h_{1}=k$. Thus, Problem 1 is also $\mathbf{NP}$-hard.
@@ 4.2. Khatri-Rao Deep Clustering
In the autoencoder-based deep clustering setting, autoencoders consist of an encoder $f^{e}_{\Theta_{\boldsymbol{\alpha}}}:\mathcal{D}\to Z$ mapping data to latent representations $Z=\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$, and a decoder $f^{d}_{\Theta_{\boldsymbol{\alpha}}}:Z\to\hat{\mathcal{D}}$, reconstructing data from their latent representations. We consider an autoencoder with $n_{l}$ layers. The $l$-th layer is parameterized by a matrix $\mathbf{W}_{l}\in\mathbb{R}^{d_{l}\times m_{l}}$, which could be relatively large. To obtain a succinct representation of the output of deep clustering algorithms, it is thus not sufficient to compress the centroid parameters, but it is also necessary to compress the autoencoder parameters $\Theta_{\boldsymbol{\alpha}}=\{\mathbf{W}_{l}\}_{l=1}^{n_{l}}$. To this aim, we draw on the connection between Khatri-Rao clustering and Hadamard decomposition (ciaperoni2024hadamard) discussed in Section 2, and we reparametrize matrices $\mathbf{W}_{l}$ as Hadamard products of $q$ factors, i.e.:
@@
where $\mathbf{A}^{l}_{i}\in\mathbb{R}^{d_{l}\times r_{i}}$ and $\mathbf{B}^{l}_{i}\in\mathbb{R}^{r_{i}\times m_{l}}$. The key intuition behind such reparameterization mirrors that of Khatri-Rao clustering. The Hadamard product of $q$ matrices of ranks $r_{1},r_{2},\dots$ and $r_{q}$ can represent matrices of rank up to $\prod_{l=1}^{q}r_{l}$ while using only $2\sum_{l=1}^{q}r_{l}$ vectors. In practice, this provides an effective compression mechanism, as also demonstrated in the context of computer vision in federated learning environments (hyeon2021fedpara), where a similar reparameterization is considered.
@@ 8. Design Choices in Khatri-Rao Clustering
Choosing the number of centroids. If the budget $b$ of vectors is not fixed and one wants to identify the number of clusters that are appropriate for the data, Khatri-Rao clustering can be combined with established techniques such as $X$-Means (Pelleg2000Xmeans) or $G$-Means (Hamerly2003Learning). Here, the number of centroids is successively increased and the current parameterization is evaluated, e.g., by using the Bayesian Information Criterion (schwarz1978estimating) or by testing if certain distributional conditions are fulfilled. In Khatri-Rao clustering, increasing the number of clusters is equivalent to either increasing the cardinality of one set of protocentroids or the number of sets of protocentroids. Moreover, for deep clustering, a Khatri-Rao variant of such algorithms as DipDECK (leiber2021dip) or DeepDPM (Ronen2022deepdpm), which also optimize the number of underlying clusters, can be designed.
@@ 9.1. Settings
Datasets. We consider two synthetic dataset generators (Blobs and Classification) from the scikit-learn Python library (scikit-learn), which allow exploring the behavior of different algorithms as a function of the number of data points, features and underlying clusters. The preliminary experiments and the scalability analysis focus on those datasets.
In addition, we consider seven real-world benchmark datasets, available through the ClustPy (leiber2023benchmarking) Python library and two synthetic datasets available through the clustbench (gagolewski2022framework) Python library. Finally, we consider the Double MNIST dataset, which is created by concatenating pairs of digits from MNIST (lecun1998gradient), yielding $100$ clusters representing the numbers 0 to $99$.
@@
In all deep clustering experiments, the parameter $w_{rec}$, which balances the clustering and reconstruction losses as per Equation (2), is set to $1$, the batch size for batch-wise gradient-based optimization is set to $512$, and we use the ADAM optimizer (Diederik2015adam) using a learning rate of $10^{-3}$ for pre-training the autoencoder and $10^{-4}$ for clustering.
@@
Metrics. In our experiments, the accuracy of a data summary is measured by the quality of clustering results. We consider the following standard metrics to evaluate clustering results that take advantage of ground truth labels: adjusted Rand index (ARI) (hubert1985comparing), unsupervised clustering accuracy (ACC) (Yang2010accuracy) and normalized mutual information (NMI) (kvalseth1987entropy). These metrics have a maximum value of $1$, which reveals a perfect match between predicted and ground-truth labels. In the case of $k$-Means clustering, we also monitor inertia, which is the objective function optimized for by standard $k$-Means and Khatri-Rao-$k$-Means. In addition, in the preliminary experiments, we monitor clustering purity, which measures the proportion of correctly assigned data points after assigning each data point to the majority ground-truth label of its cluster (manning2008introduction).
@@
Clustering algorithms. In the case of $k$-Means clustering, we compare Khatri-Rao-$k$-Means against the standard $k$-Means algorithm (implementation from scikit-learn (scikit-learn)). Additionally, we consider the two-phase naïve approach described in Section 5 with product aggregator. When not clear from context, we specify the aggregator function and the budget of vectors used to represent centroids (e.g., Khatri-Rao-$k$-Means-++ $(h_{1}+h_{2})$ denotes Khatri-Rao-$k$-Means with sum aggregator and $h_{1}+h_{2}$ protocentroids).
In the case of deep clustering, we consider DKM and IDEC (implementations from ClustPy (leiber2023benchmarking)) against their Khatri-Rao variants, Khatri-Rao DKM and Khatri-Rao IDEC. Appendix B provides implementation details for all algorithms.
@@ 9.4. Case Studies
Obtaining more succinct codebooks for color quantization. Color quantization in computer graphics enables efficient compression of certain types of images by reducing the number of colors. It allows displaying images with many colors on hardware-constrained devices that can only display a limited number of colors, usually due to memory limitations (brun2017color). A common approach casts color quantization as clustering in a three-dimensional RGB space, where pixels are points. The centroids obtained by $k$-Means form a codebook of representative colors, and each pixel is mapped to its closest centroid for color quantization.
@@
Reducing communication costs for clustering in federated learning environments. In recent years, federated learning, which allows for collaborative training of machine learning models across distributed entities, has emerged as a central area of research (wen2023survey). The Khatri-Rao clustering paradigm can benefit federated learning by reducing communication costs between entities, an important goal in federated learning research (paragliola2022definition).
Recently, garst2024federated have introduced an effective implementation of $k$-Means in a federated setting (henceforth called FkkM). As FkkM requires communicating centroids between several clients and a server, Khatri-Rao-FkkM can potentially reduce the costs of communication by considering a lightweight set of protocentroids instead of regular centroids. To extend FkkM to Khatri-Rao clustering, it suffices to replace each invocation of $k$-Means in FkkM with Khatri-Rao-$k$-Means. The resulting algorithm is referred to as Khatri-Rao-FkkM. To demonstrate the advantages given by Khatri-Rao-FkkM with respect to communication costs, we simulate a federated learning environment with $10$ clients and one server. In this experiment, we consider the benchmark FEMNIST dataset (caldas2018leaf), and we measure the inertia achieved by FkkM and Khatri-Rao-FkkM as a function of the communication costs from the server to the clients. The results of this experiment are displayed in Figure 10 and indicate that Khatri-Rao-FkkM consistently reduces inertia relative to FkkM at parity communication cost. For the smallest communication cost, the inertia incurred by FkkM is about five times larger than that incurred by Khatri-Rao-FkkM. This suggests that Khatri-Rao-FkkM can communicate clustering results of a given quality much more efficiently than FkkM.
@@ Appendix A Dataset Details
  • The MNIST dataset (lecun1998gradient), available via PyTorch (paszke2017automatic) or ClustPy (leiber2023benchmarking), is a collection of $28\times 28$ (vectorized) images of handwritten grayscale digits. We draw a stratified subsample of $25000$ images, ensuring that the original class proportions across the ten digit clusters are preserved. We rescale all pixel values by dividing by the maximum.
@@
  • The HAR dataset, available via the UCI repository (kelly2023uci) or ClustPy (leiber2023benchmarking), consists of sensor readings collected from smartphone accelerometers and gyroscopes during human activity monitoring. There are $10299$ data points. Each data point is represented as a multivariate feature vector of dimension $561$ derived from raw time-series measurements, and the dataset includes $6$ activity clusters such as walking, standing and sitting. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Olivetti Faces dataset, available via ClustPy (leiber2023benchmarking) or scikit-learn (scikit-learn) contains (vectorized) grayscale facial images of $40$ individuals, with $10$ images per subject captured under varying lighting conditions, facial expressions, and poses. Each image has resolution $64\times 64$ pixels. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The CMU Faces dataset, available via the UCI repository (kelly2023uci) or ClustPy (leiber2023benchmarking), contains $624$ grayscale facial images of $20$ persons varying their pose (up, straight, left and right) and expression (neutral, happy, sad, angry), and shown with and without sunglasses. The original images with a resolution of $30\times 32$ pixels are vectorized and each feature is standardized by subtracting its mean and dividing by its standard deviation.
@@
  • The Symbols dataset, available via ClustPy [^6] The dataset is also available at: https://www.timeseriesclassification.com/index.php (leiber2023benchmarking), consists of vectorized handwritten symbols. Each sample corresponds to a time series obtained from the drawing trajectory of a symbol. The symbols are drawn by $13$ different individuals. The dataset contains a total of $1020$ data points, each with $398$ measurements. The data points are organized into $6$ natural clusters. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The stickfigures dataset (gunnemann2014smvc), available via ClustPy (leiber2023benchmarking), consists of $900$ synthetic (vectorized) silhouette images of human stick figures generated under varying poses. Each image has a resolution of $20\times 20$ pixels. Examples of such images are given in Figure 1. Each image is provided at a fixed resolution and captures a simplified body configuration defined by joint positions and limb orientations, resulting in $9$ distinct pose-based clusters. We rescale all pixel values by dividing by the maximum.
@@
  • The optdigits dataset, available via the UCI repository (kelly2023uci) or ClustPy (leiber2023benchmarking), contains $5620$ (vectorized) handwritten digit images represented as $8\times 8$ grayscale pixel grids. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Classification dataset, sourced by scikit-learn (scikit-learn), is a synthetic dataset. While it was originally designed for benchmarking classifiers, it is also useful for evaluating clustering algorithms which simply discard class labels until the evaluation of the clustering results. Each class for classification corresponds to a cluster. Except in the experiments where we vary the number of data points, features, or clusters, the dataset consists of $5000$ samples organized into $100$ clusters. The data are generated with $10$ informative features, with no redundant or repeated features, ensuring that all dimensions contribute meaningfully to class separability. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Chameleon dataset, available via the clustbench Benchmark Suite (gagolewski2022framework), consists of $10000$ two-dimensional point clouds exhibiting complex, nonconvex cluster shapes with varying densities (DBLP:journals/computer/KarypisHK99). The dataset corresponds to the configuration shown at the bottom of Figure 4, augmented with a substantial proportion of uniformly distributed data points that do not belong to any natural cluster.
@@
  • The Soybean Large dataset, available via the UCI repository (kelly2023uci) or ClustPy (leiber2023benchmarking), is a dataset of categorical features. It consists of $562$ plant samples belonging to one of $15$ classes. For each plant, there are $35$ observed categorical attributes describing the plant. After encoding categorical attributes numerically, we standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The Blobs dataset, sourced by scikit-learn (scikit-learn), is a synthetic dataset specifically designed for controlled evaluation of clustering algorithms that consists of data points grouped around isotropic Gaussian clusters. Each cluster has standard deviation $1$. The dataset consists of $5000$ 22-dimensional data points, arranged in $100$ clusters, except in the experiments where we vary the number of data points, features and clusters. We standardize each feature by subtracting its mean and dividing by its standard deviation.
@@
  • The R15 dataset, available via the clustbench benchmark suite (gagolewski2022framework), consists of $600$ two-dimensional data points forming $15$ Gaussian clusters. The cluster centroids in R15 are not arranged on a regular grid but exhibit non-uniform inter-cluster distances. We standardize each feature by subtracting its mean and dividing by its standard deviation.